OpenStudy (anonymous):

x + y = 4 x - y = 2 solve by substitution

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

x=6 and y=-2 plz click the good answer button if u like this answer

OpenStudy (anonymous):

\[x-y=2\implies x=y+2\implies 2y+2=4\implies y=1 \implies x=3\]

OpenStudy (anonymous):

x=3 y=1

OpenStudy (anonymous):

sorry its x=3, y=1